OpenStudy (benlogic):

What is the solution of log3x + 4 4096 = 4? Can someone help me solve this?

OpenStudy (benlogic):

\[\log_{3x+4}4096=4 \]

jhonyy9 (jhonyy9):

rewrite it 4^(3x+4) = 4096 hope this will help you

jhonyy9 (jhonyy9):

ben how can you rewriting the 4096 like 4 on what power ?

OpenStudy (benlogic):

im still not sure how to do it. Could you solve it step by step maybe so I can see how you get the answer?

OpenStudy (prathamesh_m):

There's an easier way: \[4096=4^{6}\]

OpenStudy (benlogic):

im trying to figure out what x equals

OpenStudy (prathamesh_m):

So in other words: \[\log_{3x+4}4^{6}=4 \] \[3x+4=6\]

jhonyy9 (jhonyy9):

so you need to know that the function of logarithms is invers of exponential function

OpenStudy (prathamesh_m):

should be easy now

jhonyy9 (jhonyy9):

yes Parathamesh exactly

jhonyy9 (jhonyy9):

3x+4=6 and just this you need to solve it for x

OpenStudy (benlogic):

so x=2/3?

jhonyy9 (jhonyy9):

yes exactly

OpenStudy (benlogic):

that's not one of my answer choices. hold on ill show you

OpenStudy (benlogic):

the answers are: x=-1 x=0 x=4/3 x=3

jhonyy9 (jhonyy9):

4*4*4*4*4*4 = 4096

OpenStudy (prathamesh_m):

oops sorry i'm wrong

OpenStudy (prathamesh_m):

@jhonyy9 shouldn't it be : \[(3x+4)^{4}=4^{6}\]

OpenStudy (benlogic):

so id just solve that?

OpenStudy (prathamesh_m):

@benlogic u could probably just substitute the options in that equation to find the answer because expanding that would make it very complicated.

jhonyy9 (jhonyy9):

logb(y) = x => y = b^x

jhonyy9 (jhonyy9):

log_b y = x so than y = ^x

jhonyy9 (jhonyy9):

y = b^x

OpenStudy (benlogic):

im still confused, im not sure what to do

jhonyy9 (jhonyy9):

so in this case log_b y mean log_(3x+4) 4096 yes ?

jhonyy9 (jhonyy9):

3x+4 is b and 4096 mean y

OpenStudy (benlogic):

okay, I see that now. what do I do with that

OpenStudy (prathamesh_m):

From that u get the equation: \[(3x+4)^{4}=4^{6}\]

OpenStudy (prathamesh_m):

To keep it simple u can just substitute each option in that and check which one is correct.

OpenStudy (benlogic):

I think I got it!! x=4/3

jhonyy9 (jhonyy9):

so using this formula log_b y = x => y = b^x

OpenStudy (prathamesh_m):

Yes that's right

OpenStudy (benlogic):

thank, jhonny and pratha, you guys helped a lot. bye

jhonyy9 (jhonyy9):

was our pleasure

jhonyy9 (jhonyy9):

bye bye